### v1

# Deep Learning Improves Photometric Redshifts in All Regions of Color Space

Abstract
Photometric redshifts (photo-’s) are crucial for the cosmology, galaxy evolution, and transient science drivers of next-generation imaging facilities like the Euclid Mission, the Rubin Observatory, and the Nancy Grace Roman Space Telescope. Previous work has shown that image-based deep learning photo- methods produce smaller scatter than photometry-based classical machine learning (ML) methods on the Sloan Digital Sky Survey (SDSS) Main Galaxy Sample, a testbed photo- dataset. However, global assessments can obscure local trends. To explore this possibility, we used a self-organizing map (SOM) to cluster SDSS galaxies based on their colors. Deep learning methods achieve lower photo- scatter than classical ML methods for all SOM cells. The fractional reduction in scatter is roughly constant across most of color space with the exception of the most bulge-dominated and reddest cells where it is smaller in magnitude. Interestingly, classical ML photo-’s suffer from a significant color-dependent attenuation bias, where photo-’s for galaxies within a SOM cell are systematically biased towards the cell’s mean spectroscopic redshift and away from extreme values, which is not readily apparent when all objects are considered. In contrast, deep learning photo-’s suffer from very little color-dependent attenuation bias. The increased attenuation bias for classical ML photo- methods is the primary reason why they exhibit larger scatter than deep learning methods. This difference can be explained by the deep learning methods weighting redshift information from the individual pixels of a galaxy image more optimally than integrated photometry.
1 Introduction
Next-generation imaging surveys such as the Euclid mission (Laureijs et al., 2011), the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), and the Nancy Grace Roman Space Telescope (Roman; Spergel et al., 2015; Akeson et al., 2019) will obtain deep multi-band optical and near-infrared imaging over thousands of square degrees. Nearly all of the extragalactic science drivers of these surveys, such as cosmology, galaxy evolution, and transients, will require redshift estimates (Newman & Gruen, 2022). Ideally, spectroscopic redshifts (spec-’s) would be used to estimate distances, but obtaining spectra for any significant fraction of the photometrically-detected objects is completely infeasible with current facilities (e.g., Newman et al., 2015). Instead, these imaging surveys must rely on photometric redshifts (photo-’s; e.g., The LSST Dark Energy Science Collaboration et al. 2018).
Photometry-based supervised ML algorithms have produced more precise photo- estimates for imaging surveys than physically-motivated techniques, such as template fitting (which involves fitting galaxy spectral templates to a galaxy’s observed integrated photometry), when sufficient training data is available (Salvato et al., 2019). More recently, the advent of convolutional neural networks (CNNs) has enabled deep learning photo- models that can leverage pixel-level color information from the images directly (see Huertas-Company & Lanusse 2023 for a review of deep learning methods applied to galaxy surveys). These modern imaged-based deep neural networks produce the current best photo- predictions if given training samples with complete coverage of color–magnitude–redshift space, such as the Sloan Digital Sky Survey (SDSS), even with training sets that are relatively small (e.g., Pasquet et al., 2019; Hayat et al., 2021; Henghes et al., 2022; Dey et al., 2022a; Treyer et al., 2024; Ait Ouahmed et al., 2024). Despite these success, it remains unclear exactly why they perform better than classical ML methods on global metrics, which motivates this work.
Strong performance on global metrics can mask offsetting systematic errors in local regions of input parameter space. An extreme case of this effect is the trainZ photo- estimator from the LSST Photo- Data Challenge (Schmidt et al., 2020), which simply predicted the redshift distribution of the training set as the photo- probability distribution function for all objects, regardless of their photometry. By construction, the trainZ estimator achieved the ideal distribution of the probability integral transform when considered globally, beating out the other 11 widely-used photo- methods. However, it performed the worst in local regions of color space (as characterized by the conditional density estimation loss; see, e.g., Izbicki & Lee 2017). A full conditional assessment by regressing across color space (e.g., using the Cal-PIT method of Dey et al. 2022b) is beyond the scope of this paper. Rather, we simply partition color space using a self-organizing map (SOM; Kohonen 1982, 2001; see Section 4) and assess photo- performance in neighborhoods of color space.
The paper is organized as follows. Section 2 describes the SDSS dataset used in this analysis, and Section 3 reviews the photo- algorithms used. Section 4 introduces our method of partitioning the galaxy sample in color space with a SOM. Section 5 presents results on the photo- scatter and bias for individual SOM cells. Finally, Section 6 discusses the contribution of attenuation bias to photo- scatter for classical ML methods.
2 Data
Our analysis uses the SDSS (Gunn et al., 1998, 2006; York et al., 2000; Smee et al., 2013) Main Galaxy Sample (MGS; Strauss et al. 2002) from Data Release 12 (DR12; Alam et al. 2015). We started with an identical selection criteria as Pasquet et al. (2019); namely, we selected galaxies based on their dereddened -band Petrosian magnitude () and spec- (), which yielded 516,525 galaxies. We split this sample into the same training (464,873 galaxies) and test (51,652 galaxies) sets as Dey et al. (2022a). We further removed galaxies with very faint dereddened model magnitudes (, , , , or ) to ensure reliable photometry. These cuts reduced our dataset to 509,731 galaxies, with 458,740 galaxies in our training set and 50,991 galaxies in our test set.
We also used galaxy properties from the SDSS photometric catalog (Stoughton et al., 2002), such as fracDev_r (the weight of the de Vaucouleurs component in a combined de Vaucouleurs + exponential profile fit in the -band) and expAB_r (the semi-minor/semi-major axis ratio for the exponential profile fit in -band, which we note is only reliable for spirals), which we adopt as the bulge-to-total ratio and inclination, respectively. Throughout this work, we use the bulge-to-total ratio as a single parameter to summarize a galaxy’s rest-frame color (leveraging the well-known trend between morphology and rest-frame color for low- galaxies; e.g., Blanton et al. 2003), instead of the specific star formation rate, which requires highly accurate spectra that not all surveys can obtain, including SDSS.
3 Photometric Redshifts
To compare the differences between photometry-based classical ML photo- methods and image-based deep learning photo- methods, we considered two approaches in each category. For classical ML methods, we adopted the photo-’s from the SDSS DR12 catalog as estimated by Beck et al. (2016, hereafter B16) and our own photo-’s determined using a random forest. For the deep learning models, we investigated photo-’s from Dey et al. (2022a) and Pasquet et al. (2019).
3.1 SDSS DR12 Catalog
For the SDSS DR12 catalog photo-’s, B16 used local linear regression, which fits a first-order polynomial to a small number of neighbors with labels (i.e., spec-’s) in the input space (modelMag colors and cModelMag magnitude). Performing this calculation locally preserves the small-scale structure in the data and allows it to track subtle trends. B16 used 100 neighbors in their fitting procedure, though they implemented more nuanced criteria to handle under-sampled regions of color–magnitude space and to iteratively filter redshift outliers.
B16 trained their model using 1,976,978 galaxies with spec-’s from the entire SDSS DR12 spectroscopic sample—specifically, the MGS, the luminous red galaxy sample (Eisenstein et al., 2001), and the Baryon Oscillation Spectroscopic Survey (BOSS; Dawson et al., 2013)—plus additional spectroscopy from other surveys, such as DEEP2 (Newman et al., 2013), VIPERS (Garilli et al., 2014; Guzzo et al., 2014), and zCOSMOS (Lilly et al., 2009). B16 estimated photo-’s for the 208,474,076 galaxies of the SDSS primary photometric catalog, thus their aims were broader than those of the other approaches we consider, which only used MGS galaxies for training and testing. That being said, MGS galaxies dominate the information content within the color–magnitude–redshift space covered by MGS galaxies, which is further prioritized by the local nature of their regression method.
3.2 Random Forest
Inspired by previous random forest photo- efforts (e.g., Carliles et al., 2010; Carrasco Kind & Brunner, 2013; Zhou et al., 2019), we trained a random forest (Breiman, 2001) for photo- estimation using the RandomForestRegressor method from the Python package Scikit-learn (Pedregosa et al., 2011) as a classical ML alternative to the local linear regression method of B16. Random forest regression is a simple supervised learning algorithm consisting of a “forest” of decision trees that are each trained to predict a value with the final result of the algorithm being the average of each tree’s predicted value. To train the algorithm, we used the training set described in Section 2. We tested providing the random forest with multiple different sets of input features, starting with the features B16 used to train their network, as discussed in Section 3.1. To test the impact of input feature choices, we ran random forest models with several combinations of input features:
1.
modelMag colors and cModelMag -band magnitude (our fiducial set up);
2.
modelMag colors, cModelMag -band magnitude, modelMag color uncertainties, and cModelMag -band magnitude uncertainty;
3.
modelMag colors, modelMag -band magnitude, the axis ratio for the exponential model (expAB_r), and the bulge-to-total ratio (fracDev_r); and
4.
modelMag colors, modelMag -band magnitude, modelMag color uncertainties, modelMag -band magnitude uncertainty, the axis ratio for the exponential model (expAB_r), and the bulge-to-total ratio (fracDev_r).
Ultimately, we found only minor differences in the photo-’s by adding more features, so we chose model 1 (the same features as B16) as our fiducial random forest photo- model. However, we adopt the B16 photo-’s as our fiducial classical ML photo- method because it produced qualitatively similar results to the random forest approach, but the B16 photo-’s are widely used and are a standard comparison sample for photo- analyses.
3.3 Deep Learning Models
Despite their architectural differences, a variety of modern deep learning models—including CNNs (Pasquet et al., 2019), semi-supervised representation learning (Hayat et al., 2021), and deep capsule networks (Dey et al., 2022a)—have all significantly out-performed classical ML photo- methods (namely B16) on the SDSS MGS. In this work, we considered photo-’s from Pasquet et al. (2019) and Dey et al. (2022a), both of which were trained using the dataset described in Section 2. We adopt the latter as our fiducial deep learning photo- model, though we found very similar results with the former.
3.3.1 Capsule Networks
Hinton et al. (2011) proposed the idea of capsule networks as an alternative to CNNs, where the artificial neurons of a capsule network are organized into groups (capsules), producing the vector analog to the scalar neurons in CNNs such that they are viewpoint equivariant (as opposed to the viewpoint invariance of CNNs) and therefore could be more efficient to train. However, actually training capsule networks proved challenging until the introduction of the dynamic routing algorithm by Sabour et al. (2017) that dramatically improved the viability of their training. Soon after, Katebi et al. (2019) became the first work to apply capsule networks to galaxy images for galaxy morphology prediction.
Dey et al. (2022a) utilized a deep capsule network for their encapzulate model (compared to the shallow network proposed by Sabour et al. 2017) developed by Rajasegaran et al. (2019). Its architecture consists of three sub-networks: a deep capsule network for encoding the galaxy images to a low dimensional embedding space, a class-independent decoder network for morphological prediction that acts as a regularizer (using disk/spheroid labels from Galaxy Zoo 1; Lintott et al. 2011), and a redshift prediction network.
3.3.2 Global Photo-z Performance
Figure 1 shows versus for the B16 and encapzulate photo-’s of SDSS MGS test set galaxies. Compared to the B16 photo-’s, the encapzulate photo-’s have a smaller scatter as measured using the normalized median absolute deviation:
where . In the case of photo-’s, we define as (, ). The encapzulate photo-’s also have a significantly lower fraction of catastrophic outliers (as defined by , where ) and a smaller global bias () than the B16 photo-’s. These are generic results of comparing classical ML and deep learning photo- methods for the SDSS dataset, though it is not understood how the deep learning models leverage additional redshift information from the images.
Furthermore, there is a slight but perceptible trend of the B16 photo-’s being biased towards the mean redshift of the dataset, such that they are overestimated at low and underestimated at high , which is a common bias for ML methods called attenuation bias (see Section 6.1). Below, we compare classical ML and deep learning photo-’s for sub-samples in small neighborhoods in color space to understand how local trends impact global metrics.
4 Partitioning Color Space with a Self-Organizing Map
To study the impact of attenuation bias in local neighborhoods of color space, we partition color space using a SOM. A SOM is an unsupervised dimensionality reduction algorithm that works by grouping together objects with similar properties into discrete cells, with objects in a cell having properties that are most similar to other objects in the same cell and neighboring cells usually containing galaxies that are broadly similar. Since SOMs produce easily interpretable results with clear visualizations, they have been used for photo- estimation (e.g., Carrasco Kind & Brunner 2014; Myles et al. 2021; Wright et al. 2025) and characterizing the coverage of color space by objects with spec-’s (e.g., Masters et al. 2015, 2017; Stanford et al. 2021; Gruen et al. 2023; McCullough et al. 2024).
For our implementation, we used the minisom python package (Vettigli, 2018) to produce a two-dimensional SOM with 10 12 cells using the , , , and colors of all objects in our sample (i.e., including objects in both the training and test sets). Figure 2 shows the SOM color-coded by each cell’s mean spec- (left panel) and mean bulge-to-total ratio (B/T; right panel). Roughly speaking, mean spec- increases from left-to-right and B/T increases from bottom-to-top.
As representatives of the two main galaxy populations, we chose Cell A (8, 9), which contains red elliptical galaxies (high B/T), and Cell B (1, 3), which contains blue disks (low B/T). The other two cells of interest show more unusual scenarios. Cell C (3, 4) has an intermediate B/T and is located in the transition region between low and high B/T. It contains a diverse collection of objects, including red disk galaxies (Masters et al., 2010), post-starburst galaxies (Dressler & Gunn, 1983), and galaxies with active galactic nuclei, determined by examining the SDSS and DESI-LS images and SDSS spectra of galaxies in this cell. Finally, Cell D (0, 0) contains very low- galaxies (), a regime where the B16 algorithm suffered from a large fraction of catastrophic outliers (Geha et al., 2017) but deep learning methods performed well (e.g., Wu et al., 2022; Darragh-Ford et al., 2023).
5 Results
In this section, we compare the scatter and bias of classical ML and deep learning photo-’s in localized regions of color space (i.e., in individual SOM cells) and investigate trends of these quantities with the mean bulge-to-total ratio in each cell (B/T).
5.1 Photo-z Scatter in Individual SOM Cells
Figure 3 shows the ratio between the values of the B16 and encapzulate photo-’s for individual SOM cells (gray points) as a function of the cell’s B/T. For all of the SOM cells, the for the B16 photo-’s is at least as large as that of encapzulate with most of the cells having an value – times larger. The binned median trend (solid blue line) shows that there is a mostly flat but slightly decreasing trend from B/T with an ratio of about –, which is similar to the global ratio of 1.5. The highest B/T cells typically have smaller from encapzulate, though a few individual cells (like Cell A) approach parity.
We found similar results when using the random forest photo-’s in lieu of the B16 photo-’s and when using the Pasquet et al. (2019) photo-’s instead of the encapzulate photo-’s. This suggests that the deep learning methods are harnessing additional redshift information from the spatially-resolved images that purely photometry-based ML methods cannot access, particularly for SOM cells with low and intermediate B/T.
5.2 Slope of – in Individual SOM Cells
Figure 4 shows images (first column) from the Dark Energy Spectroscopic Instrument Legacy Surveys (DESI-LS; Dey et al. 2019) and the – plots for three photo- methods (B16, encapzulate, and random forest; second, third, and fourth columns, respectively) with the rows corresponding to Cells A–D (as described in Section 4). In the – comparisons, the black solid line denotes the one-to-one line, and points outside of the black dashed lines are considered outliers using the criterion . The red lines show the best-fit lines for each SOM cell determined with least-trimmed-squares regression using the LtsFit package (Cappellari et al., 2013). This algorithm works by iteratively clipping outliers in the regression fit (not to be confused with photo- outliers in the – diagram). For performing the regression, we adopted the of the objects in a given cell (for a given photo- method) as the estimated uncertainty. Although uncertainties should be negligible, we assumed an uncertainty of to allow for catastrophic spec- failures such as blends or shredded objects.
For Cells A–C, the best-fit lines for the classical ML photo- methods (B16 and random forest) are significantly flatter than for the deep learning photo- method (encapzulate), which instead closely track the one-to-one line in all panels. In spite of the large fraction of outliers, the best-fit line for the B16 photo-’s in Cell D is close to the one-to-one line because many objects were trimmed in the regression.
We determined the angle of the best-fit – line by taking the arc-tangent of the slope for each cell, measured counterclockwise from the positive direction, which we denote by . A perfect one-to-one line would correspond to . The left panel of Figure 5 shows the resulting values for each SOM cell as a function of B/T for the B16 (blue) and encapzulate (orange) photo-’s. The right panel shows the marginal distributions of ; in both panels, the gray line indicates the ideal () angle. Due to the attenuation bias, the B16 photo-’s have systematically lower values () than the encapzulate photo-’s (), which are close to ideal. We note that the values of the B16 photo-’s approach those of the encapzulate photo-’s at the very highest B/T values, mimicking the trend seen in Figure 3.
The values from our random forest photo-’s and the Pasquet et al. (2019) deep learning photo-’s are qualitatively similar to those from B16 and encapzulate, respectively, so we have omitted those results from Figure 5 for clarity. This suggests that our findings are likely a generic result when comparing image-based deep learning methods to classical ML photo- methods.
6 Discussion
6.1 The Impacts of Attenuation Bias Versus Random Scatter
As we have shown, photometry-based classical ML photo-’s exhibit a factor of larger scatter than image-based deep learning photo-’s both when considered globally (see Figure 1) and in local regions of color space (see Figure 3). Although the global – trends for both classical ML and deep learning photo-’s align with the one-to-one line (with values of 42∘ and 44∘, respectively), classical ML photo-’s show much flatter slopes in – plots when restricted to objects in a local neighborhood of color space (e.g., an individual SOM cell) compared to deep learning photo-’s, which remain close to the ideal one-to-one line. For classical ML photo-’s, the flatter – trends combine to form a global – trend that tracks the one-to-one line but with more scatter than if there were no bias. Figure 6 illustrates schematically how the – relations for individual color cells combine for classical ML and deep learning photo-’s.
The flatter local – slopes obtained from classic ML methods are a manifestation of a common phenomenon that affects both linear regression and ML methods, known as attenuation bias or regression dilution. This effect causes output values to be biased towards the mean (i.e., with predictions that are too high when values are lowest and too low when values are highest) due to the impact of measurement errors in the independent variable/feature values (e.g., Fuller 1987; Akritas & Bershady 1996; Loredo 2004; Kelly 2007; Ting 2024). In this section, we assess the relative contributions of attenuation bias and random scatter to the overall scatter within individual SOM cells.
We expect that the overall scatter (within a SOM cell) can be decomposed into two components whose sum in quadrature should be roughly equal to the overall scatter (if standard deviation were used as the metric rather than , this relationship would be exact). First, we compute the component of the scatter due to attenuation bias: = (, ), where is the value of the cell’s best fit – relation at each object’s . Second, we calculate the scatter of the observed values in a given cell around that cell’s best fit – relation: = (, ).
Figure 7 shows and its two components ( and ) for each cell as a function of B/T for both the B16 (left column) and encapzulate (center column) photo-’s. The right column of Figure 7 shows / (middle row) and / (bottom row) for each cell and photo- method. For the B16 photo-’s, and provide roughly similar contributions to the overall scatter; however, for the encapzulate photo-’s, has a much smaller contribution to the overall scatter than . In contrast, the values are only a slightly lower for encapzulate than for B16. We also obtain qualitatively similar results to encapzulate using the Pasquet et al. (2019) photo-’s, and similar results to B16 using our random forest photo-’s. Hence, we conclude that the overall lower photo- errors obtained with deep learning methods applied to SDSS can primarily be attributed to their much lower level of attenuation bias, rather than to decreases in random scatter.
6.2 Deep Learning Leverages Pixels to Mitigate Attenuation Bias
Ting (2024) investigated strategies to mitigate attenuation bias by leveraging many features that are correlated because they result from the same underlying physical process (see Sections 2.4.2 and 2.5.2 of that work). They presented a simple scenario where many noisy features (in our case, the individual pixels in a galaxy image) are linearly correlated with a physical property (here, redshift). Often those features are combined to increase the signal-to-noise ratio (e.g., integrated photometry) before predicting the physical property. If the relative weightings of features in the combination step is perfect, then using individual features for prediction does not mitigate the attenuation bias. However, if the relative weightings are imperfect, then predicting from individual features will reduce the attenuation bias. We have verified these predictions with Monte Carlo models for a toy scenario.
In our case, deep learning photo- models utilize the spectral energy distributions (SEDs) of many individual pixels per object that all share the same redshift (ignoring unassociated blends, which are very rare at the depths of the SDSS MGS; Dawson et al. 2016) but have uncorrelated noise. The integrated photometry used for the B16 photo-’s is based upon model-fitting photometry which should (if the model is accurate) weight pixels proportionally to an object’s flux. If the photometry model matches the true light distribution of the galaxy perfectly and all pixels’ SEDs are equally sensitive to redshift, then it would be an ideal weighting of the pixels and hence would be expected to yield comparable attenuation bias to using all the pixels as independent features for prediction.
Instead, the attenuation bias in photo-’s is reduced for predictions using the actual pixel values compared to integrated photometry, particularly for bluer or less bulge-dominated galaxies. This suggests that the photometry models for these galaxies are less accurate and/or that some pixels are more informative about redshift than their relative flux implies. Indeed, we would expect that for galaxies with significant spiral structure or irregularity, the simple composite exponential/De Vaucoleurs models used for SDSS photometry would not capture all of the variations in intensity amongst pixels, making the first of these scenarios plausible for later-type galaxies but of little import for early-type galaxies that tend to have much smoother morphologies. Furthermore, late-type galaxies exhibit much stronger gradients in colors and stellar population properties than early-type galaxies do (de Vaucouleurs, 1961); the colors of the older, redder populations at the centers of these galaxies are more strongly correlated with redshift than the colors of their outer portions (Gladders & Yee, 2000).
As a result, both of these explanations for our results are very plausible: later-type galaxies have larger deviations of pixel fluxes from the models used for integrated photometry and also have larger variations in pixel sensitivity to redshift than would be expected from their relative brightnesses in the SDSS photometry models alone. This can cause redshift predictions from integrated photometry to suffer from greater attenuation bias than predictions from the raw pixel values, especially for galaxies with smaller bulge-to-total ratios, as we have found. Our results are not sufficient, however, to indicate which of these two possible effects is dominant. One potential approach for disentangling this puzzle would be to make redshift predictions from photometry integrated within multiple circular annuli (which should tend to separate components of different colors; e.g., Saxena et al. 2024) and compare to the attenuation bias and overall – relations when predicting redshift from raw pixel values.
6.3 Outlook
While a full investigation of how deep learning photo- methods are able to mitigate the local attenuation bias that affects classical methods is beyond the scope of this work, a better understanding of how deep learning models are able to utilize the information available in images may provide key insights for designing more efficient photo- strategies, such as new data augmentation procedures for contrastive learning approaches that utilize noise augmentations. In particular, these efforts will have important implications for developing deep learning photo- estimation algorithms for Euclid (Euclid Collaboration et al., 2025), LSST (Merz et al., 2025), and Roman (Khederlarian et al. in prep.).